# Convergence analysis of Levenberg-Marquardt method with Singular Scaling for non-zero residue nonlinear least-squares problems [^1] This work is supported by FAPESC (Fundação de Amparo à Pesquisa e Inovação do Estado de Santa Catarina) [grant number 2023TR000360]

### v1

@@ 5 Global Convergence
where $M_{3}=L_{J}^{2}+M_{0}$. We conclude that
$-\nabla\phi(x_{k})^{T}d_{k}\geq\lambda_{k}\|d_{k}\|^{2}\geq\frac{\lambda_{k}}{M_{3}^{2}}\|\nabla\phi(x_{k})\|^{2}=\frac{1}{M_{3}^{2}}\|\nabla\phi(x_{k})\|^{3}$
if we take $\lambda_{k}=\|\nabla\phi(x_{k})\|$. Moreover,
$\|d_{k}\|=\|(J_{k}^{T}J_{k}+\lambda_{k}I)^{-1}\nabla\phi(x_{k})\|$ $\leq\|(J_{k}^{T}J_{k}+\lambda_{k}I)^{-1}\|\|\nabla\phi(x_{k})\|\leq 1,$
which shows that our “safeguard directions” are also gradient-related.
@@
Now that it is proven that Algorithm 1 generates a sequence of directions $\{d_{k}\}$ that is gradient-related, using [4, Proposition 1.2.1], global convergence can be established.
Theorem 4.
@@
Let $\{x_{k}\}$ be a sequence generated by Algorithm 1. Then, every limit point $\hat{x}$ of $\{x_{k}\}$ is such that $\nabla\phi(\hat{x})=0$.
Proof.
@@
Let $K_{1}=\{k\in\mathbb{N}\mid\|J(x_{k}+d_{k})^{T}F(x_{k}+d_{k})\|\leq\vartheta\|J(x_{k})^{T}F(x_{k})\|\}$. If $K_{1}$ is infinite, it follows that $\|J(x_{k})^{T}F(x_{k})\|\rightarrow 0$, and therefore any limit point $\hat{x}$ of $\{x_{k}\}$ is such that $J(\hat{x})^{T}F(\hat{x})=0$, hence $\nabla\phi(\hat{x})=0$. Otherwise, if $K_{1}$ is finite, let us assume, without loss of generality, that $\|J(x_{k}+d_{k})^{T}F(x_{k}+d_{k})\|>\vartheta\|J(x_{k})^{T}F(x_{k})\|$, for each $k$, such that the step size is chosen to satisfy the Armijo condition. Since, by Proposition 1 and Remark 5, the directions of Algorithm 1 are gradient-related, it follows from [4, Proposition 1.2.1] that any limit point $\hat{x}$ of $\{x_{k}\}$ is a stationary point of $\phi(x)$. ∎∎
@@
The next result establishes the connection between global and local convergence by showing that, under certain conditions, $\alpha_{k}=1$ for all $k$ sufficiently large and then Algorithm 1 becomes the “pure” LMMSS and the results of Section 4 are applicable.
Theorem 5.
@@
Let $\{x_{k}\}$ be generated by Algorithm 1, using $x_{0}\in\mathbb{R}^{n}$ as the initial point. Suppose that Assumptions 1-3 and Assumption 4, with $r=1$, hold at the level set ${\cal L}_{0}=\{x\in\mathbb{R}^{n}\mid\phi(x)\leq\phi(x_{0})\}$. Moreover, assume that ${\cal L}_{0}$ is compact. Then, for every $k$ sufficiently large $\alpha_{k}=1$ and the sequence $\{\textup{dist}(x_{k},X^{*})\}$ converges to 0 quadratically.
Proof.
Note that by the definition of Algorithm 1 $\{\phi(x_{k})\}$ is decreasing. Starting from $x_{0}$, we conclude that the sequence $\{x_{k}\}$ remains in ${\cal L}_{0}$, a compact set. Therefore, there exists a convergent subsequence, as per Theorem 4, where every limit point of this subsequence is stationary. By removing this convergent subsequence, we can repeat this process successively for the remaining sequence, which remains in a compact set, thereby concluding that $\{x_{k}\}$ is the disjoint and enumerable union of convergent subsequences, each of which has a stationary point as limit.
@@
Thus, for every $k$ sufficiently large, the norm of the gradient at $x_{k}$ will be sufficiently small. In other words, there is a positive integer $k_{0}$ such that
@@
$\leq\frac{\tilde{C}L_{3}\|J(x_{k})^{T}F(x_{k})\|}{\omega^{3}}\|J(x_{k})^{T}F(x_{k})\|$ (58)
@@
proving that $\alpha_{k}=1$ for each $k\geq k_{0}$. To obtain the quadratic convergence of $\{\textup{dist}(x_{k},X^{*})\}_{k\geq k_{0}}$ to 0, one may apply Theorem 3 when we have a diminishing rank or from Theorem 2, when $rank(J(x))=rank(J(x^{*}))\geq 1$. ∎∎
Remark 6.
We observe that in the case of zero residue a previous result [13, Theorem 3.1] assumed that the limit point $x^{*}$ is such that $F(x^{*})=0$. Also, in the case of unconstrained optimization, in order to prove a result similar to Theorem 5, the required assumption is that the Hessian at $x^{*}$ is positive definite; however, such assumption would imply that $x^{*}$ is an isolated stationary point. As we want to address the case of non-zero residue and possibly non-isolated stationary points, we considered the boundness of the level set instead of these other two conditions.
Remark 7.
In the scenario of diminishing rank, if Assumption 4 is satisfied with $10<r<$ 1, it is necessary to adjust the parameter $\lambda_{k}$ of Algorithm 1 to $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert^{r}$. Furthermore, when this hypothesis holds, the above results follow with minor changes in some constants, and we can show there exists $k_{0}$ such that $\alpha_{k}=1,\forall k\geq k_{0}$ and that $\textup{dist}(x_{k},X^{*})$ goes to zero superlinearly.
@@ 6 Illustrative examples
The purpose of this section is to illustrate our convergence results by presenting some examples of non-zero residue nonlinear least-squares problems. In this section we use the shorthand notation (A1) – (A4) for Assumptions 14, respectively.
Example 1. Let us start with an example from [2, Example 5.1] with non-isolated minimizers (stationary-points). Consider the residual function
@@
has rank $1$ everywhere except for the origin. Moreover, as discussed in [2], Assumptions (A2), (A3) and (A4) with $r=1$ hold as well.
By considering LMMSS with the scaling matrix $L=[-1\quad 1]$, Assumption (A1) holds everywhere except for the line $x_{2}=-x_{1}$.
Hence, we are under conditions to apply the results of Section 4.1. As expected from Theorem 2, in Table 1, we can observe the local quadratic convergence of $\text{dist}(x_{k},X^{*})$ to zero starting from two distinct starting points in a neighborhood of $x^{*}=(0,\sqrt{5})^{T}\in X^{*}$.
@@
This example is also useful to illustrate the effect of using $L^{T}L\neq I$ and the globalization mechanisms of Algorithm 1.
Figure 1a depicts the trajectory of LMM ($L^{T}L=I$) and the one of LMMSS (with $L=[-1\quad 1]$) both starting from $x_{0}=(2,4)^{T}$. As we can see from the figure, while LMM dives perpendicularly to the set $X^{*}$ (observe the level curves in the figure), LMMSS trajectory is parallel to the null space of $L$, and they end up at different stationary points. As exposed in [5] for parameter identification problems, an advantage of using Algorithm 1 is that in a given problem with non-isolated solutions, we can choose $L$ that enforces a specific property. In this geometric example, by using $L=[-1,1]$, we aim for the method to prioritize stationary points where $\|L(x-x_{0})\|=|x_{2}-x_{1}-2|$ is as small as possible. In [5], the authors employed $L$ as discrete versions of derivative operators in order to promote smoothness in the obtained solutions for parameter identification problems in heat conduction. Moreover, Figure 1b enforces this aspect by showing that the presence of a well-thought choice of $L$ may conduce the iterations faster and in a more accurate manner than LMM, while it stimulates a desired property.
@@
The last aspect we shall discuss with this example is about the completeness condition and the global convergence mechanisms of Algorithm 1. As mentioned before, (A1) holds everywhere except for the line $x_{2}=-x_{1}$. LMMSS also tries to move parallel to the null space of $L$. Figure 2a illustrates what happens if we apply LMMSS without the safeguard Step 4, using $x_{0}=(-1,3)^{T}$. The displacements $x_{k}-x_{0}$ are in ${\cal N}(L)$, but the hyperplane $x_{0}+{\cal N}(L)$ does not intersect $X^{*}$. As a result, LMMSS minimizes $\phi(x)$ in $x_{0}+{\cal N}(L)$ approximating a point $\hat{x}$ where $\nabla\phi(\hat{x})\perp{\cal N}(L)$ so that the directions $d_{k}$ become orthogonal to $-\nabla\phi(x_{k})$ loosing the descent property and the algorithm crashes not satisfying the Armijo condition. Notice that $\hat{x}$ is such that $\hat{x}_{2}=-\hat{x}_{1}$ and (A1) fails. On the other hand, we use Algorithm 1 with the safeguard step, Step 4 corrects this behavior and when $x_{k}$ approaches $\hat{x}$ the classic LMM direction is employed instead; see Figure 2b. This allows the algorithm to move on and reach $X^{*}$.
@@
Example 2. Now we consider an example to illustrate the “diminishing rank” scenario. The residual function is given by
@@
which leads to $\phi(x)$ having an isolated global minimizer at $(-1,0)^{T}$ and a non-isolated set of local minimizers $\{x\in\mathbb{R}^{2}\mid x_{1}=0\}$ that we shall use as $X^{*}$ and, in this case, $\textup{dist}(x,X^{*})=|x_{1}|$.
This example was also analyzed in [2] where it was verified that assumptions (A2), (A3) and (A4) with $r=1$ hold in a neighborhood of $x^{*}=(0,2)^{T}$.
For assumption (A1), we point out that the null space of
$J(x)=\left[\begin{array}[]{cc}3x_{1}^{2}-x_{2}&-x_{1}\\ 3x_{1}^{2}+x_{2}&x_{1}\end{array}\right]$
is $\{0\}$ wherever $x_{1}\neq 0$ and it is $\text{span}\{(0,1)^{T}\}$ for $x\in X^{*}\setminus\{0\}$. Hence, it is clear that, regardless the choice of $L$, (A1) holds everywhere except for $X^{*}$.
In Table 2 and Figure 3 we report the behavior of LMM and LMMSS with $L=[-1\quad 1]$ starting in the neighborhood of $x^{*}=(0,2)^{T}$. From Table 2 the quadratic convergence rate is apparent for both methods. In Figure 3 we highlight the different limit points: while the classic LMM moves perpendicularly to $X^{*}$, LMMSS moves parallel to ${\cal N}(L)$, leading in this case to a stationary point with smaller $x_{2}$ coordinate.
@@
Example 3. Let us consider $F(x)=(x_{1}^{2},x_{2}^{2},x_{1}+x_{2},1),$ the residual function. The corresponding least-squares function $\phi(x)$ has a unique global minimizer at $x^{*}=(0,0)^{T}$, which is the only stationary point of $\phi(x)$, $X^{*}=\{x^{*}\}$. The rank of
$J(x)=\left[\begin{array}[]{cc}2x_{1}&0\\ 0&2x_{2}\\ 1&1\\ 0&0\end{array}\right]$
is 2 for $x\neq x^{*}$ and it is 1 at $x^{*}$. Therefore the rank diminishes as we approach $X^{*}$ and (A1) holds everywhere, regardless $L$, except at $x^{*}$. Also, since $\|J(x)-J(y)\|\leq 2\|x-y\|$, (A2) is verified.
The distinguished feature of this example is that the error bound condition (A3) does not hold for every point in a neighborhood of $x^{*}$. In fact, by considering the sequence $x_{k}=(1/k,-1/k)^{T}$ it is not hard to show that
$\lim_{k\to\infty}\frac{\|\nabla\phi(x_{k})\|}{\textup{dist}(x_{k},X^{*})}=\lim_{k\to\infty}\frac{\|\nabla\phi(x_{k})\|}{\|x_{k}\|}=0,$
and thus (A3) cannot hold.
Nevertheless, for any $x\in\text{span}\{(1,1)^{T}\}={\cal N}(J(x^{*}))^{\perp}$, from the mean value theorem, we have
$\|\nabla\phi(x)-\nabla\phi(0)\|=\|\nabla^{2}\phi(\xi)x\|\geq 2\|x\|,$
where we used the fact that $(1,1)^{T}$ is an eigenvector of $\nabla^{2}\phi(\xi)$ and 2 is a lower bound for the corresponding eigenvalue. Therefore, in this example, starting at $x_{0}\in{\cal N}(J(x^{*}))^{\perp}$ and using $L$ such that ${\cal N}(L)={\cal N}(J(x^{*}))^{\perp}$, we force the sequence $\{x_{k}\}$ generated by LMMSS to stay in a subspace where (A3) holds. Also, with this choice of $L$, (A1) is satisfied everywhere, and since we have
$\|J(x)^{T}F(0)\|=0\leq K\|x-0\|^{2},$
(A4) with $r=1$ is fulfilled and quadratic convergence is achieved, as depicted in Table 3.
Table 3: Distances of $x_{k}$ to $x^{*}$ produced by LMMSS until the stopping criterion $\|J_{k}^{T}F_{k}\|<10^{-10}$ is achieved for Example 3 with $L=[-1\quad 1]$, for different starting points.
$k$ $x_{0}^{(1)}=(3,3)^{T}$ $x_{0}^{(2)}=(-2,-2)^{T}$
1 $2.0097$ $1.2571$
2 $8.0542\times 10^{-1}$ $3.8494\times 10^{-1}$
3 $1.5845\times 10^{-1}$ $2.4840\times 10^{-2}$
4 $1.9403\times 10^{-3}$ $7.6586\times 10^{-6}$
5 $3.6524\times 10^{-9}$ $2.6331\times 10^{-16}$
6 $7.4734\times 10^{-17}$
It is worth mentioning that even though (A3) does not hold for every $x$ in a neighborhood of $x^{*}$ in this example, a weaker Hölderian error bound condition [1] does hold. Precisely $\|\nabla\phi(x)\|\geq\|x\|^{3}$ or, in other words, $\nabla\phi$ is Hölder metrically subregular of order 1/3 (see [1, Definition 1]).

### v2

@@ 5 Global convergence
where $M_{3}=L_{J}^{2}+\bar{\lambda}$. We conclude that
$-\nabla\phi(x_{k})^{T}d_{k}\geq\lambda_{k}\|d_{k}\|^{2}\geq\frac{\lambda_{k}}{M_{3}^{2}}\|\nabla\phi(x_{k})\|^{2}=\frac{1}{M_{3}^{2}}\|\nabla\phi(x_{k})\|^{2+r}$
if we take $\lambda_{k}=\|\nabla\phi(x_{k})\|^{r}$, or $-\nabla\phi(x_{k})^{T}d_{k}\geq\frac{\underline{\lambda}}{M_{3}^{2}}\|\nabla\phi(x_{k})\|^{2}$, if $\lambda_{k}\geq\underline{\lambda}>0$. Moreover,
$\|d_{k}\|=\|(J_{k}^{T}J_{k}+\lambda_{k}I)^{-1}\nabla\phi(x_{k})\|$ $\leq\|(J_{k}^{T}J_{k}+\lambda_{k}I)^{-1}\|\|\nabla\phi(x_{k})\|$
$\leq\frac{1}{\lambda_{k}}\|\nabla\phi(x_{k})\|=\|\nabla\phi(x_{k})\|^{1-r},$
if $\lambda_{k}=\|\nabla\phi(x_{k})\|^{r}$, or $\|d_{k}\|\leq\|\nabla\phi(x_{k})\|/\underline{\lambda}$ if $\lambda_{k}\geq\underline{\lambda}>0$. In both cases the right rand sides are bounded in the convergent subsequence $\{x_{k}\}_{k\in{\cal K}}$.
@@
Now that it is proven that Algorithm 1 generates a sequence of directions $\{d_{k}\}$ that is gradient-related, using (Bertsekas2016, , Proposition 1.2.1), global convergence can be established.
Theorem 6
@@
Given $x_{0}\in\mathbb{R}^{n}$, let $\{x_{k}\}$ be a sequence generated by Algorithm 1 with either $\lambda_{k}=\|J_{k}^{T}F_{k}\|^{r},r\in]0,1[$ or $0<\underline{\lambda}\leq\lambda_{k}\leq\bar{\lambda}$. Then, every limit point $\hat{x}$ of $\{x_{k}\}$ is such that $\nabla\phi(\hat{x})=0$.
Proof
@@
Let $K_{1}=\{k\in\mathbb{N}\mid\|J(x_{k}+d_{k})^{T}F(x_{k}+d_{k})\|\leq\vartheta\|J(x_{k})^{T}F(x_{k})\|\}$. If $K_{1}$ is infinite, it follows that $\|J(x_{k})^{T}F(x_{k})\|\rightarrow 0$, and therefore any limit point $\hat{x}$ of $\{x_{k}\}$ is such that $J(\hat{x})^{T}F(\hat{x})=0$, hence $\nabla\phi(\hat{x})=0$. Otherwise, if $K_{1}$ is finite, let us assume, without loss of generality, that $\|J(x_{k}+d_{k})^{T}F(x_{k}+d_{k})\|>\vartheta\|J(x_{k})^{T}F(x_{k})\|$, for each $k$, such that the step size is chosen to satisfy the Armijo condition. Since the directions of Algorithm 1 are gradient-related (due to Proposition 1 and discussion afterwards), it follows from (Bertsekas2016, , Proposition 1.2.1) that any limit point $\hat{x}$ of $\{x_{k}\}$ is a stationary point of $\phi(x)$. ∎
@@
The next result establishes the connection between global and local convergence by showing that, under certain conditions, $\alpha_{k}=1$ for all $k$ sufficiently large and then Algorithm 1 becomes the “pure” LMMSS and the results of Section 4 are applicable. This is the only result in this section which requires Assumption 4.
Theorem 7
@@
Let $\{x_{k}\}$ be generated by Algorithm 1, using $x_{0}\in\mathbb{R}^{n}$ as the initial point and $\lambda_{k}=\|J_{k}^{T}F_{k}\|$. Suppose that Assumption 2, Assumption 4 with $r=1$, and condition (4) hold at the level set ${\cal L}_{0}=\{x\in\mathbb{R}^{n}\mid\phi(x)\leq\phi(x_{0})\}$. Moreover, assume that ${\cal L}_{0}$ is compact.
Then, for every $k$ sufficiently large, $\alpha_{k}=1$, and the sequence $\{\textup{dist}(x_{k},X^{*})\}$ converges quadratically to zero.
Proof
@@
Note that by the definition of Algorithm 1 $\{\phi(x_{k})\}$ is decreasing. Starting from $x_{0}$, we conclude that the sequence $\{x_{k}\}$ remains in ${\cal L}_{0}$, a compact set. By Theorem 6, every limit point $\hat{x}$ of $\{x_{k}\}$ is stationary, thus for $k$ sufficiently large, the norm of the gradient at $x_{k}$ will be sufficiently small. In other words, there is a positive integer $k_{0}$ such that
@@
$\leq\frac{\tilde{C}L_{3}\|J(x_{k})^{T}F(x_{k})\|}{\omega^{3}}\|J(x_{k})^{T}F(x_{k})\|$
@@
proving that $\alpha_{k}=1$ for each $k\geq k_{0}$. To obtain the quadratic convergence of $\{\textup{dist}(x_{k},X^{*})\}_{k\geq k_{0}}$ to 0, one may now apply Theorem 3 (or Theorem 4) with $x^{*}=\hat{x}$ and $x_{0}=x_{k_{0}}$. ∎
Remark 7
We observe that, in the case of zero residual, a previous result (Yamashita, , Theorem 3.1) assumed that the limit point $x^{*}$ is such that $F(x^{*})=0$. Also, in the case of unconstrained optimization, in order to prove a result similar to Theorem 7 the required assumption is that the Hessian at $x^{*}$ is positive definite; however, such assumption would imply that $x^{*}$ is an isolated stationary point. As we want to address the case of nonzero residual and possibly nonisolated stationary points, we considered the boundness of the level set instead of these other two conditions.
Remark 8
In the scenario of diminishing rank, if Assumption 4 is satisfied with $10<r<$ 1, it is necessary to adjust the parameter $\lambda_{k}$ of Algorithm 1 to $\lambda_{k}=\left\lVert J_{k}^{T}F_{k}\right\rVert^{r}$. Furthermore, when this hypothesis holds, Theorem 7 follows with minor changes in some constants, and we can show there exists $k_{0}$ such that $\alpha_{k}=1,\forall k\geq k_{0}$ and that $\textup{dist}(x_{k},X^{*})$ goes to zero superlinearly.
@@ 6 Illustrative examples
The purpose of this section is to illustrate our convergence results by presenting some examples of nonzero residual nonlinear least-squares problems. In this section we use the shorthand notation (A1) – (A4) for Assumptions 14, respectively. The examples came from reference (behling2019local, , Section 5), where (A1)–(A4) were already demonstrated/verified and such verification will not be repeated here.
Example 1. (Constant rank and (A4)) Let us start with an example from (behling2019local, , Example 5.1) with nonisolated minimizers (stationary-points). Consider the residual function
@@
has rank $1$ everywhere except for the origin. Moreover, as discussed in behling2019local , (A1), (A2) and (A4) with $r=1$ hold as well. It is worth to point out that (A4) holds with $C=0$ because in this example $\|J(x)^{T}F(z)\|=0$.
By considering LMMSS with the scaling matrix $L=[-1\quad 1]$, condition (4) holds everywhere except for the line $x_{2}=-x_{1}$.
Hence, if we consider an initial point in a neighborhood of $x^{*}=(0,\sqrt{5})^{T}\in X^{*}$, we are under conditions to apply the results of Section 4.1. As expected from Theorem 3, in Table 1, we can observe the local quadratic convergence of $\text{dist}(x_{k},X^{*})$ to zero, starting LMMSS from two distinct starting points in a neighborhood of $(0,\sqrt{5})^{T}$.
@@
This example is also useful to illustrate the effect of using $L^{T}L\neq I$ and the globalization mechanisms of Algorithm 1. Figure 2a depicts the trajectory of LMM ($L^{T}L=I$) and the one of LMMSS (with $L=[-1\quad 1]$) both starting from $x_{0}=(2,4)^{T}$. As we can see from the figure, while LMM dives perpendicularly to the set $X^{*}$ (observe the level curves in the figure), LMMSS trajectory seems parallel to the null space of $L$, and they end up at different stationary points but with the same residual norm. Figure 2b plots the gradient norm through the iterations of LMM and LMMSS.
@@
The last aspect we shall discuss with this example is about the completeness condition (4) and the global convergence mechanisms of Algorithm 1. As mentioned before, condition  (4) holds everywhere except for the line $x_{2}=-x_{1}$. LMMSS also tries to move parallel to the null space of $L$. Figure 3a illustrates what happens if we apply LMMSS without the safeguard Step 4, using $x_{0}=(-1,3)^{T}$. The displacements $x_{k}-x_{0}$ are in ${\cal N}(L)$, but the hyperplane $x_{0}+{\cal N}(L)$ does not intersect $X^{*}$. As a result, LMMSS minimizes $\phi(x)$ in $x_{0}+{\cal N}(L)$, approximating a point $\hat{x}$ where $\nabla\phi(\hat{x})\perp{\cal N}(L)$ so that the directions $d_{k}$ become orthogonal to $-\nabla\phi(x_{k})$ loosing the descent property and the algorithm crashes not satisfying the Armijo condition. Notice that $\hat{x}$ is such that $\hat{x}_{2}=-\hat{x}_{1}$ and condition (4) fails. On the other hand, if we use Algorithm 1 with the safeguard step, Step 4 corrects this behaviour and, when $x_{k}$ approaches $\hat{x}$, the classic LMM direction is employed instead; see Figure 3b. This allows the algorithm to move on and reach $X^{*}$.
@@
Example 2. (Diminishing rank and (A4)) Now we consider an example to illustrate the “diminishing rank” scenario. The residual function is given by
@@
which leads to $\phi(x)$ having an isolated global minimizer at $(-1,0)^{T}$ and a nonisolated set of local minimizers $\{x\in\mathbb{R}^{2}\mid x_{1}=0\}$ that we shall use as $X^{*}$ and, in this case, $\textup{dist}(x,X^{*})=|x_{1}|$.
This example was also analyzed in (behling2019local, , Example 5.2) where it was verified that assumptions (A1), (A2) and (A4) with $r=1$ hold in a neighborhood of $x^{*}=(0,2)^{T}$. Differently from the previous example where $J(x)^{T}F(z)=0$, here (A4) holds because $\|J(x)^{T}F(z)\|=6x_{1}^{2}=6\textup{dist}(x,X^{*})^{2}$.
For condition (4), we point out that the null space of
$J(x)=\left[\begin{array}[]{cc}3x_{1}^{2}-x_{2}&\quad-x_{1}\\ 3x_{1}^{2}+x_{2}&\quad x_{1}\end{array}\right]$
is $\{0\}$ wherever $x_{1}\neq 0$ and it is $\text{span}\{(0,1)^{T}\}$ for $x\in X^{*}\setminus\{0\}$. Hence, it is clear that, regardless the choice of $L$, condition (4) holds everywhere except for $X^{*}$.
In Table 2 and Figure 4 we report the behaviour of LMM and LMMSS, with $L=[-1\quad 1]$, starting in the neighbourhood of $x^{*}=(0,2)^{T}$. From Table 2, the quadratic convergence rate is apparent for both methods. In Figure 4 we highlight the different limit points: while the classic LMM moves almost perpendicularly to $X^{*}$, LMMSS moves parallel to ${\cal N}(L)$, leading in this case to a stationary point with smaller $x_{2}$ coordinate.
@@
Example 3. (Diminishing rank and (A3)) The residual function
$F(x)=(\frac{1}{9}\cos x_{1}-x_{2}\sin x_{1},\ \ \frac{1}{9}\sin x_{1}+x_{2}\cos x_{1})^{T}$
has a set of nonisolated minimizers $X^{*}=\{x\in\mathbb{R}^{2}\mid x_{2}=0\}$, thus $\textup{dist}(x,X^{*})=|x_{2}|$. In (behling2019local, , Example 5.3) it was verified that for $x^{*}=(\pi,0)^{T}$, and $\delta=0.1$, (A1) is satisfied, as well as (A2) with $\omega=1$ (because $\|\nabla\phi(x)\|=|x_{2}|$).
The Jacobian, given by
$J(x)=\left[\begin{array}[]{cc}-x_{2}\cos x_{1}-\frac{1}{9}\sin x_{1}&\quad-\sin x_{1}\\ \\ -x_{2}\sin x_{1}+\frac{1}{9}\cos x_{1}&\quad\cos x_{1}\end{array}\right]$
has full rank for $x_{2}\neq 0$ and rank 1 for $x_{2}=0$. Thus, the rank of $J(x_{k})$ decreases as $x_{k}$ approaches $X^{*}$ from the outside. Moreover, condition (4) is valid in $B(x^{*},\delta)\setminus X^{*}$.
Unfortunately, (A4) does not hold in this example. But (A3) does, because
$\|J(x)^{T}F(z)\|$ $=\frac{1}{18}\sqrt{81\sin^{2}(x_{1}-z_{1})+(9x_{2}\cos(x_{1}-z_{1})+\sin(x_{1}-z_{1}))^{2}}$
$\leq\frac{\sqrt{91}}{81}\|x-z\|,$
for $x\in B(x^{*},\delta)$ and $z\in X^{*}\cap B(x^{*},\delta)$. Hence, $\sigma=\sqrt{91}/81\approx 0.118$.
Other constants relevant for the analysis are given approximately by $L_{0}=1.1156$, $L_{4}=0.5612$, $\gamma=0.8949$, $(\sigma_{\min}^{*})^{2}=1.0123$. We observe that $\sigma<(\sigma_{\min}^{*})^{2}$, and for $\theta=1.0315$, we ensure $(\sigma_{\min}^{*})^{2}>\theta\sigma$. In this case, $\bar{\delta}=0.1=\delta$, and assumptions required by Lemma 6 hold.
By considering $L=[1\quad 0]$, $\|L\|=1$ and from (42) with $\kappa=1.0001$ we obtain $c_{1}=2.6172$. Then, since $\bar{\sigma}=0.1367>\sigma$, we can follow Lemma 14 with $\varepsilon<=0.0016$ and $\eta\in]0.8617,1[$.
Setting $\eta=0.9$, makes $\varepsilon=0.0016$ in (43) and starting from $x_{0}=(\pi,0.001)^{T}$, and using $\lambda_{k}=\sigma+L_{4}\|x_{k}-\bar{x}_{k}\|$, Theorem 5 applies, and we expect linear convergence, as can be verified in Table 3.
$k$ $\textup{dist}(x_{k},X^{*})$
0 0.001
1 0.00010413
2 1.0889e-05
3 1.1392e-06
4 1.1919e-07
5 1.247e-08
6 1.3046e-09
7 1.365e-10
8 1.4281e-11
Table 3: Iterations of LMMSS for Example 3, initial point $x_{0}=(\pi,0.001)^{T}$.